# Axionlike particle-assisted supercooling chiral phase transition in QCD: Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenarios

### v2

@@ III.2 The case with $\mu_{B}\neq 0$
The first term defined as $\Omega_{\mu_{B}}^{(1)}$, up to a term constant in $M$, goes like $\sim\frac{N_{c}N_{f}}{12\pi^{2}}(\frac{\mu_{B}}{3})^{2}M^{2}$ for a large $\mu_{B}\gg M$, which is the Debye mass screening. The second term $\Omega_{\mu_{B}}^{(2)}\sim\frac{-N_{c}N_{f}}{16\pi^{2}}M^{4}\ln\frac{\left[(\frac{\mu_{B}}{3})+\sqrt{(\frac{\mu_{B}}{3})^{2}-M^{2}}\right]^{2}}{M^{2}}$ takes a scale anomaly form of the CW type and, in an extreme limit $\mu_{B}\gg M$, precisely cancels the original one at $\mu_{B}=0$ in Eq.(9). A potential barrier is generated at moderately large $\mu_{B}$ by creation of the Fermi surface, acting as a repulsive force, which is dictated by $\Omega_{\mu_{B}}^{(1)}$. This triggers the first order phase transition at $T=0$.
At finite $T$, the cancellation for the original scale anomaly form of the CW-type becomes incomplete, hence it will survive during the thermal phase transition, as long as the Debye mass screening by $\mu_{B}$ efficiently works so as to make the dynamical chiral symmetry breaking milder. Thus, a large enough $\mu_{B}$ can control the size of the soft-scale breaking, and at finite $T$, can pull the type of the first order phase transition back to the CW-type.
Figure 1: The variation of the thermodynamic potential at low $T$ ($T=20$ MeV) with respect to $\mu_{B}$, in unit of ${\rm MeV}^{4}$, as a function of $S=\frac{M-m}{2(g_{s}+g_{d})}$ at $\theta=0$ in unit of ${\rm MeV}^{3}$, in the MFA-NJL description monitoring two-flavor QCD at physical point (for details, see the text).
@@
Figure 1 demonstrates how the CW-type potential can be realized by finite $\mu_{B}$ at low enough $T$. The appropriate size of $\mu_{B}$ would thus be $\sim 1100$ MeV, which is much greater than the typical QCD chemical potential size of ${\cal O}(300)$ MeV. Requiting this size of $\mu_{B}$ is also consistent with the analysis in the literature  Shao:2024dxt ; Shao:2024ygm , to achieve sizable nano hertz GW production from quark-meson models. Note that in Fig. 1, the soft-scale breaking effect $\propto S^{2}$ contribute negatively at around $S\sim 0$. Here, to fix the model parameters, we have referred to empirical hadron observables in the isospin symmetric limit at the vacuum Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs : the pion mass $m_{\pi}=140.2\,\mathrm{MeV}$, the pion decay constant $f_{\pi}=92.6\,\mathrm{MeV}$, and quark condensate $\langle\bar{u}u\rangle=\langle\bar{d}d\rangle=(-241.5\,\mathrm{MeV})^{3}$. Then the model parameters can be determined as Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ,
@@
In Fig. 2 we show the thermal evolution of the CW-type chiral phase transition with $\mu_{B}=1100$ MeV fixed. The CW supercooling can indeed be achieved in this setup. The stochastic GW background can then be produced and the predicted spectra could be similar to those discussed in Shao:2024dxt ; Shao:2024ygm , where a two-flavor quark meson model with a large $\mu_{B}$ such as $\mu_{B}\sim 1$ GeV has been shown to generate nano hertz GW signals, without clarifying the limit to the CW-type chiral phase transition. As inspired also in the literature Shao:2024dxt ; Shao:2024ygm , the Affleck-Dine baryogenesis Affleck:1984fy ; Linde:1985gh ; Dine:2003ax could produce such a large $\mu_{B}$ in the QCD phase transition epoch.
@@ IV ALP assisted CW-type chiral phase transition in QCD
We start with considering a new scalar field ($\Phi$) coupled to $M$, like $\Phi^{2}M^{2}$ of the portal form, and allow $\Phi$ to develop the vacuum expectation value and generate an extra soft-scale breaking mass term, to cancel the dynamical chiral symmetry breaking at low-enough $T$. Suppose that this vacuum expectation value becomes nonzero in the chiral symmetric phase at high $T$, and almost vanishes at the vacuum with $T=0$. This setup ensures that the extra soft-scale breaking effect is activated to trigger the cosmological CW-type phase transition at $T=T_{c}$, and below $T_{n}$, it freezes out to make the QCD-chiral broken phase normal, so that the present-day QCD hadron physics is intact.
@@
The decoupling of extra soft-scale breaking can be realized, say, if the vacuum expectation value of $\Phi$ scales with $\mu_{5}$, the chiral chemical potential, so the $\Phi$ field acts like an ALP coupled to the axial charge, as has been discussed in Zhang:2025kbu . This decoupling is then encoded with the freezing out of the QCD sphaleron, which would happen in the epoch of the confinement-deconfinement phase transition, which takes place seemingly simultaneously with the chiral phase transition. As long as the QCD sphaleron is active in hot QCD plasma, the chiral chemical potential $\mu_{5}$ along with a local CP-odd domain is created Manton:1983nd ; Klinkhamer:1984di , so that the QCD vacuum characterized by the strong CP phase $\theta$ and its fluctuation (in the spatial-homogeneous direction) gets significantly sizable Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe within the QCD time scale McLerran:1990de ; Moore:1997im ; Moore:1999fs ; Bodeker:1999gx : the sphaleron transition rate is not suppressed by the thermal effect in contrast to the instanton’s one Moore:1997im ; Moore:1999fs ; Bodeker:1999gx . The time fluctuation $\partial_{t}\theta(t)$, to be referred to as the chiral chemical potential, $\mu_{5}$ Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe ; Andrianov:2012hq ; Andrianov:2012dj , will be significant as well when the non-conservation law of the $U(1)$ axial symmetry is addressed.
@@
This is definitely a negative mass-squared term ensured by the gauge covariance for the external axialvector field $A_{\mu}=\mu_{5}\delta_{\mu}^{0}$, which corresponds to the repulsive force from the vector interaction against the scalar probes charged under the corresponding gauge.
Thus, combined with the quartic $\lambda|\Phi|^{4}$ term, $\Phi$ can get nonzero vacuum expectation value, $v_{\Phi}=\frac{2\mu_{5}}{\sqrt{\lambda}}$. We assume that the $\Phi$ field is settled at the vacuum expectation value at a high enough $T$, even when $T\gtrsim T_{c}$, where the QCD sphaleron actively generates $\mu_{5}$. At $T\sim T_{c}$, the Yukawa coupling term in Eq.(16), when put in the current MFA-NJL framework, modifies the thermodynamic potential to deform the type of the chiral phase transition. It makes a shift of the scalar ($S$) and pseudoscalar ($P$) mean fields terms only in the loop correction parts, by $(y\,\phi_{s})$ and $(y\,\phi_{p})$, respectively. Without loss of generality, we take $\langle\phi_{s}\rangle=v_{\Phi}$ and $\langle\phi_{p}\rangle=0$. Keeping $\theta=0$, hence $P=0$, we then find that up to thermal correction parts, the vacuum contribution in Eq.(10) looks like modified as
@@
$\sim\frac{N_{f}N_{c}}{8\pi^{2}}(yv_{\Phi})^{2}=\frac{N_{f}N_{c}}{2\pi^{2}}\frac{y^{2}\mu_{5}^{2}}{\lambda}\,$, (21)
up to the logarithmic factor of ${\cal O}(1)$. Thus, the total ${\cal M}^{2}$ term goes like
$\Omega_{M^{2}}$ $\sim\left[\frac{1}{2(g_{s}+g_{d})}-\frac{N_{f}N_{c}}{4\pi^{2}}\Lambda^{2}+\frac{N_{f}N_{c}}{8\pi^{2}}(yv_{\Phi})^{2}\right]{\cal S}^{2}+\cdots\,$,
$\sim\frac{N_{f}N_{c}}{8\pi^{2}}\left[(yv_{\Phi})^{2}-(475\,{\rm MeV})^{2}\right]{\cal S}^{2}+\cdots\,$, (22)
@@
where ellipses stand for thermal corrections terms at $T\gtrsim T_{c}$, and we have quoted values of the model parameters referenced in Eq.(15). As clearly seen from Eq.(22), the soft-scale breaking mass term can be canceled when $(yv_{\Phi})\sim 475$ MeV, so that the chiral phase transition almost goes like the CW-type supercooling.

### v4

@@ III.2 The case with $\mu_{B}\neq 0$
The first term defined as $\Omega_{\mu_{B}}^{(1)}$, up to a term constant in $M$, goes like $\sim\frac{N_{c}N_{f}}{12\pi^{2}}(\frac{\mu_{B}}{3})^{2}M^{2}$ for a large $\mu_{B}\gg M$, which is the Debye mass screening. The second term $\Omega_{\mu_{B}}^{(2)}\sim\frac{-N_{c}N_{f}}{16\pi^{2}}M^{4}\ln\frac{\left[(\frac{\mu_{B}}{3})+\sqrt{(\frac{\mu_{B}}{3})^{2}-M^{2}}\right]^{2}}{M^{2}}$ takes a scale anomaly form of the CW type and, in an extreme limit $\mu_{B}\gg M$, precisely cancels the original one at $\mu_{B}=0$ in Eq.(9). A potential barrier is generated at a moderately large $\mu_{B}$ by the creation of the Fermi surface, acting as a repulsive force, dictated by $\Omega_{\mu_{B}}^{(1)}$. This triggers the first order phase transition at $T=0$.
At finite $T$, the cancellation for the original scale anomaly form of the CW-type becomes incomplete, hence it will survive during the thermal phase transition, as long as the Debye mass screening by $\mu_{B}$ efficiently works so as to make the dynamical chiral symmetry breaking milder. Thus, a large enough $\mu_{B}$ can control the size of the soft-scale breaking. Still at finite $T$, thus the first order phase transition can be seen because of the residual CW-type scale anomaly effect.
Figure 1: The variation of the thermodynamic potential with respect to $\mu_{B}$ at low $T$ ($T=20$ MeV), as a function of $S=\frac{M-m}{2(g_{s}+g_{d})}$ at $\theta=0$, in the NJL-MF description monitoring two-flavor QCD at physical point (for details, see the text).
@@
Figure 1 demonstrates how the CW-type potential can be realized by finite $\mu_{B}$ at low enough $T$. The desired size of $\mu_{B}$ would thus be $\sim 1100$ MeV #4#4#4When $\mu_{B}\sim 1100$ MeV, we have the quark chemical potential $\mu_{q}=\mu_{B}/3\sim 360$ MeV. One might think that this $\mu_{q}$, which is close to the cutoff $\Lambda\simeq 590$ MeV (as in Eq.(15)), would lead to poor accuracy and a cutoff artifact in the results on the chiral phase transition. However, it is not the case: the dynamical quark mass $M=2(g_{s}+g+d)|S|\sim 364$ MeV, which is read off from Eq.(15) and the nontrivial vacuum place in Fig. 2. From these values, we estimate the Fermi momentum $p_{F}=\sqrt{\mu_{q}^{2}-M^{2}}$ to be $\sim 47$ MeV. This is the scale that plays the role of the effective cutoff for the medium effect at $T\sim 0$, above which the higher momentum contributions are exponentially suppressed. At finite $T$, the suppression is still exponential: for momenta $p\gg\mu_{q},T$, the integrand for the medium part of $\Omega$ in Eq.(6), $\Omega(T,\mu)-\Omega(T,0)$, scales as $e^{-(p-\mu_{q})/T}$, making contributions from $p\sim\Lambda$ almost negligible for $T$ of our interest ($\sim 100$ MeV). Thus, the medium correction to the chiral phase transition is utterly free and decoupled from the cutoff dependence, as long as the Fermi momentum $\Lambda$, as in the present case. . Requiring this size of $\mu_{B}$ is also consistent with the analysis in the literature  Shao:2024dxt ; Shao:2024ygm . Note that in Fig. 1, the soft-scale breaking effect $\propto S^{2}$ contributes negatively around $S\sim 0$. Here, to fix the model parameters, we have referred to empirical hadron observables in the isospin symmetric limit at vacuum Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs : the pion mass $m_{\pi}=140.2\,\mathrm{MeV}$, the pion decay constant $f_{\pi}=92.6\,\mathrm{MeV}$, and quark condensate $\langle\bar{u}u\rangle=\langle\bar{d}d\rangle=(-241.5\,\mathrm{MeV})^{3}$. Then the model parameters can be determined as Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ,
@@
In Fig. 2 we show the thermal evolution of the CW-type chiral phase transition with $\mu_{B}=1100$ MeV fixed. The CW-type supercooling can indeed be achieved in this setup. The stochastic GW background can then be produced and the predicted spectra could be similar to those discussed in Shao:2024dxt ; Shao:2024ygm , where a two-flavor quark meson model with a large $\mu_{B}$ such as $\mu_{B}\sim 1$ GeV has been shown to generate nano hertz GW signals, without clarifying the limit to the CW-type chiral phase transition. As inspired also in the literature Shao:2024dxt ; Shao:2024ygm , the Affleck-Dine baryogenesis Affleck:1984fy ; Linde:1985gh ; Dine:2003ax could produce such a large $\mu_{B}$ in the QCD phase transition epoch.
@@ IV ALP assisted CW-type chiral phase transition in QCD
IV.1 Cancellation of soft-scale breaking mass term in QCD plasma
@@
We start with considering a new scalar field ($\Phi$) coupled to $M$, like $\Phi^{2}M^{2}$ of the portal form, and allow $\Phi$ to develop the vacuum expectation value and generate an extra soft-scale breaking mass term, to cancel the dynamical chiral symmetry breaking at low-enough $T$. Suppose that this vacuum expectation value becomes nonzero in the chiral symmetric phase at high $T$, and almost vanishes at vacuum with $T=0$. This setup ensures that the extra soft-scale breaking effect is activated to trigger the cosmological CW-type phase transition at $T=T_{c}$, and below $T_{n}$, it freezes out to make the QCD-chiral broken phase normal, so that the present-day QCD hadron physics is intact.
@@
The decoupling of extra soft-scale breaking can be realized, say, if the vacuum expectation value of $\Phi$ scales with $\mu_{5}$, the chiral chemical potential, so the $\Phi$ field acts like an ALP coupled to the axial charge, as has been discussed in Zhang:2025kbu . This decoupling is then encoded with the freezing out of the QCD sphaleron, which would happen in the epoch of the confinement-deconfinement phase transition, which takes place seemingly simultaneously with the chiral phase transition. As long as the QCD sphaleron is active in hot QCD plasma, the chiral chemical potential $\mu_{5}$ along with a local CP-odd domain is created Manton:1983nd ; Klinkhamer:1984di , so that the QCD vacuum, characterized by the strong CP phase $\theta$, and the vacuum fluctuation (in the spatial-homogeneous direction) become significantly sizable Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe within the QCD time scale McLerran:1990de ; Moore:1997im ; Moore:1999fs ; Bodeker:1999gx : the sphaleron transition rate is not suppressed by the thermal effect in contrast to the instanton’s one Moore:1997im ; Moore:1999fs ; Bodeker:1999gx . The time fluctuation $\partial_{t}\theta(t)$, to be referred to as the chiral chemical potential, $\mu_{5}$ Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe ; Andrianov:2012hq ; Andrianov:2012dj , will be significant as well when the non-conservation law of the $U(1)$ axial symmetry is addressed.
This $\mu_{5}$ arises as the static solution for the quasi thermal equilibrium system McLerran:1990de , and its thermal evolution around the QCD phase transition involves a nonperturbative decoupling of the QCD sphaleron due to the confinement. which makes $\mu_{5}$ frozen out as well. Therefore, we will take $\mu_{5}$ to be constant in $T$ to be of ${\cal O}(100)$ MeV, as has conventionally been applied in the $\mu_{5}$ physics in the chiral phase transition #5#5#5The constant $\mu_{5}$ can be understood as follows: The chiral chemical potential $\mu_{5}$ in terms of thermodynamics is related to the axial number density $n_{5}$ and the axial susceptibility $\chi_{5}$ as $\chi_{5}=\frac{\partial n_{5}}{\partial\mu_{5}}$. In the thermal equilibrium, the nonconservation law of $n_{5}$ yields $n_{5}\sim\Gamma_{\rm sph}/\Gamma_{\rm chi}$, where $\Gamma_{\rm sph}$ is the QCD sphaleron-transition rate and $\Gamma_{\rm chi}$ denotes the chirality-flip rate $\sim N_{c}^{2}T^{3}/M^{2}(T)$ with the dynamical quark mass $M(T)$, which, e.g., arises via the gluon - quark elastic scattering, $q+g\leftrightarrow q+g$. Therefore, the chiral chemical potential thermally evolves like $\frac{d\mu_{5}}{dT}=\frac{dn_{5}}{dT}\frac{1}{\chi_{5}}\sim\frac{d}{dT}(\frac{T\cdot\Gamma_{\rm sph}}{M^{2}(T)})\frac{1}{\chi_{5}}$. Around the chiral criticality ($T\sim T_{c}$), $(1/M(T))$ peaks at $T\simeq T_{c}$. Since the $U(1)$ axial and $SU(2)$ chiral restorations are contaminated, and $n_{5}$ acts as the indicator for the $U(1)$ axial breaking strength – i.e,. $n_{5}\to\infty$ when the $U(1)$ axial symmetry is restored –, $n_{5}$ also exhibits a peak structure around $T\sim T_{c}$. This is a pseudocriticality when the theory is at the physical point, which follows a power law reflecting the remnant of the universality class (e.g. $O(4)$) around $T\sim T_{c}$. The QCD sphaleron rate $\Gamma_{\rm sph}$ exhibits a Boltzmann suppression $\sim\frac{N_{c}^{2}-1}{N_{c}^{2}}\cdot(\alpha_{s}N_{c})^{5}\cdot T^{4}e^{-\Lambda_{\rm QCD}/T}$ McLerran:1990de ; Moore:2010jd due to the mass gap generation (confinement). Thus, $\frac{d\mu_{5}}{dT}\sim 0$ around $T\sim T_{c}$ (for $T>T_{c}$). When the theory undergoes (the CW-type) ultra supercooling, as in the present model, below $T<T_{c}$, $\chi_{5}$ and $M$ will not drastically change with $T$ to be almost saturated respectively to the vacuum values $\sim m_{\eta^{\prime}}^{2}$ and $\sim(\langle-\bar{q}{q}\rangle_{T=0})^{1/3}$, where $m_{\eta^{\prime}}$ is the QCD $\eta^{\prime}$ mass associated with the $U(1)$ axial anomaly at the vacuum. On the other hand, $\Gamma_{\rm sph}$ still keeps Boltzmann damped when $T$ cools down til the nucleation/percolation temperature $T_{n}\ll T_{c}$. Therefore, we can conclude that $\frac{d\mu_{5}}{dT}\sim 0$ even in the case of the supercooled chiral phase transition. .
@@
This is definitely a negative mass squared term ensured by the gauge covariance for the external axialvector field $A_{\mu}=\mu_{5}\delta_{\mu}^{0}$, which corresponds to the repulsive force from the vector interaction against the scalar probes charged under the corresponding gauge.
Thus, combined with the quartic term $\lambda|\Phi|^{4}$, $\Phi$ can get nonzero vacuum expectation value, $v_{\Phi}=\frac{2\mu_{5}}{\sqrt{\lambda}}$. We assume that the $\Phi$ field settles at the vacuum expectation value at a high enough $T$, even when $T\gtrsim T_{c}$, where the QCD sphaleron actively generates $\mu_{5}$. At $T\sim T_{c}$, the Yukawa coupling term in Eq.(16), when applied to the current NJL-MF framework, modifies the thermodynamic potential to deform the type of the chiral phase transition. It makes a shift of the mean field terms of the scalar ($S$) and pseudoscalar ($P$) only in the loop correction parts, by $(y\,\phi_{s})$ and $(y\,\phi_{p})$, respectively. Without loss of generality, we take $\langle\phi_{s}\rangle=v_{\Phi}$ and $\langle\phi_{p}\rangle=0$. Keeping $\theta=0$, hence $P=0$, we then find that up to the thermal correction parts, the vacuum contribution in Eq.(10) looks like modified as
@@
$\sim\frac{3N_{f}N_{c}}{8\pi^{2}}(yv_{\Phi})^{2}=\frac{3N_{f}N_{c}}{2\pi^{2}}\frac{y^{2}\mu_{5}^{2}}{\lambda}\,$, (21)
up to the logarithmic factor of ${\cal O}(1)$. Thus, the total ${\cal S}^{2}$ term goes like
$\Omega_{{\cal S}^{2}}$ $\sim\left[\frac{1}{2(g_{s}+g_{d})}-\frac{N_{f}N_{c}}{4\pi^{2}}\Lambda^{2}+\frac{3N_{f}N_{c}}{8\pi^{2}}(yv_{\Phi})^{2}\right]{\cal S}^{2}+\cdots\,$,
$\sim\frac{3N_{f}N_{c}}{8\pi^{2}}\left[(yv_{\Phi})^{2}-(274\,{\rm MeV})^{2}\right]{\cal S}^{2}+\cdots\,$, (22)
@@
where ellipses represent thermal correction terms at $T\gtrsim T_{c}$, and we have quoted values of the model parameters referenced in Eq.(15). As clearly seen from Eq.(22), the soft-scale breaking mass term can be canceled when $(yv_{\Phi})\sim 274$ MeV.
The cancellation in Eq.(22) is thus dedicate and actually the resultant CW-type scale anomaly takes a wrong sign for generation of the first order phase transition at this point. To make it more complete, we may incorporate the meson fluctuation contribution, beyond the MFA level. Although the meson loop corrections should generically be subleading and decoupled for ${\cal S}/\Lambda\sim 1$ because the meson mass is typically given as $\sim\sqrt{\lambda_{\rm meson}}{\cal S}\ll\Lambda$ with $\lambda_{\rm meson}\gg 1$, they can play a crucial role around the criticality (at $T\sim T_{c}$). In particular, when the negative mass square of the chiral order parameter $\cal S$} is turned off as in the present scenario, the pion-loop can give a significant CW-type scale anomaly of the form $\sim+{\cal S}^{4}\ln{\cal S}$ with the right sign. This term can be a trigger for dynamical chiral symmetry breaking to determine the curvature and barrier structure around the true vacuum where the net $m_{\pi}\sim 0$, even away from the origin of the potential (${\cal S}\sim 0$#6#6#6This feature is characteristic of the present scenario, in contrast to the conventional one with a large negative mass squared at the “tree-level” of the meson sector, which dominates because all the field-dependent meson masses, including the pion mass, become heavy enough due to the large mass square, and the loop corrections are highly suppressed, when the curvature and/or barrier structure of the potential around the true vacuum is determined.. In what follows, we will more explicitly see this pion-loop trigger.
The present NJL-MFA dynamically generates the meson sector Lagrangian, which takes the form
${\cal L}_{\rm meson}=\frac{Z}{2}(\partial_{\mu}\sigma)^{2}+\frac{Z}{2}(\partial_{\mu}\pi^{a})^{2}-C_{0}\cdot\sigma-\frac{1}{2}\widetilde{m}_{0}^{2}(\sigma^{2}+(\pi^{a})^{2})-\frac{\widetilde{\lambda}_{0}(\sigma,\pi^{a})}{4}(\sigma^{2}+(\pi^{a})^{2})^{2}\,$, (23)
with
$Z$ $\sim\frac{N_{c}}{8\pi^{2}}\ln\frac{\Lambda^{2}}{{\cal S}^{2}}\,$,
$\widetilde{C}_{0}$ $\sim-(yv_{\Phi})\frac{N_{c}}{\pi^{2}}\Lambda^{2}\,$,
$\widetilde{m}^{2}_{0}$ $\sim\frac{3N_{c}}{2\pi^{2}}\left[(yv_{\Phi})^{2}-(274\,{\rm MeV})^{2}\right]\,$,
$\widetilde{\lambda}_{0}(\sigma,\pi^{a})$ $\sim\frac{N_{c}}{2\pi^{2}}\ln\frac{\Lambda^{2}}{(\sigma^{2}+(\pi^{a})^{2})}\,.$ (24)
Here we have focused only on the $(\sigma,\pi^{a=1,2,3})\approx(\bar{q}q,\bar{q}i\gamma_{5}\tau^{a}q)$ meson sector for the chiral $U(2)_{L}\times U(2)_{R}$ partners (because they are the lowest spectra) and have taken the massless quark limit $m=0$, for simplicity. This meson-sector Lagrangian can be interpreted as an effective theory evaluated at the renormalization scale $\mu={\cal S}(=\langle\sigma\rangle)$, with the compositeness condition: $Z\to 0$ and $\widetilde{\lambda}_{0}\to 0$ when $\mu\to\Lambda$. The canonical normalization of the meson fields is achieved by the redefinition $\sigma\to\frac{\sigma}{\sqrt{Z}}$ and $\pi^{a}\to\frac{\pi^{a}}{\sqrt{Z}}$, so that we have
${\cal L}_{\rm meson}\to\frac{1}{2}(\partial_{\mu}\sigma)^{2}+\frac{1}{2}(\partial_{\mu}\pi^{a})^{2}-C\cdot\sigma-\frac{1}{2}{m}_{0}^{2}(\sigma^{2}+(\pi^{a})^{2})-\frac{\lambda_{0}(\sigma,\pi^{a})}{4}(\sigma^{2}+(\pi^{a})^{2})^{2}\,$, (25)
with
$C$ $\sim-\frac{2\sqrt{2}}{\pi}\sqrt{\frac{N_{c}}{\ln\frac{\Lambda^{2}}{{\cal S}^{2}}}}(yv_{\Phi})\Lambda^{2}\,$,
${m}^{2}_{0}$ $\sim\frac{12}{\ln\frac{\Lambda^{2}}{{\cal S}^{2}}}\left[(yv_{\Phi})^{2}-(274\,{\rm MeV})^{2}\right]\,$,
${\lambda}_{0}(\sigma,\pi^{a})$ $\sim\frac{32\pi^{2}}{N_{c}[\ln\frac{\Lambda^{2}}{{\cal S}^{2}}]^{2}}\ln\frac{\Lambda^{2}}{(\sigma^{2}+(\pi^{a})^{2})}\,.$ (26)
Based on the meson-sector Lagrangian in Eq.(26), we compute the one-loop effective potential for the vacuum expectation value $\langle\sigma\rangle\equiv\sigma_{0}$ by integrating out the meson fluctuations up to one-loop level. For the large cutoff $\Lambda\gg\sigma_{0}$, the effective potential $V_{\sigma_{0}}^{\rm eff}$ is thus computed as
$V_{\sigma_{0}}^{\rm eff}=C\cdot\sigma_{0}+\frac{M^{2}(\sigma_{0})}{2}\sigma_{0}^{2}+\frac{\lambda(\sigma_{0})}{4}\sigma_{0}^{4}\,$, (27)
with
$M^{2}(\sigma_{0})$ $\sim m_{0}^{2}+\frac{\lambda_{0}(\sigma_{0},0)}{48\pi^{2}}\Lambda^{2}\,$,
$\lambda(\sigma_{0})$ $\sim\lambda_{0}(\sigma_{0},0)+\frac{3\lambda_{0}^{2}(\sigma_{0},0)}{16\pi^{2}}\ln\frac{\sigma_{0}^{2}}{\Lambda^{2}}\,.$ (28)
The meson loop-corrected potential parameters are further evaluated as
$M^{2}(\sigma_{0})$ $\sim\frac{12}{\ln\frac{\Lambda^{2}}{{\cal S}^{2}}}\left[(yv_{\Phi})^{2}-(274\,{\rm MeV})^{2}+\frac{\Lambda^{2}}{18N_{c}}\left(\frac{\ln\frac{\Lambda^{2}}{\sigma^{2}_{0}}}{\ln\frac{\Lambda^{2}}{{\cal S}^{2}}}\right)\right]\,$,
$\lambda(\sigma_{0})$ $\sim\frac{32\pi^{2}\ln\frac{\Lambda^{2}}{\sigma_{0}^{2}}}{N_{c}[\ln\frac{\Lambda^{2}}{{\cal S}^{2}}]^{2}}\left[1-\frac{6}{N_{c}}\left(\frac{\ln\frac{\Lambda^{2}}{\sigma^{2}_{0}}}{\ln\frac{\Lambda^{2}}{{\cal S}^{2}}}\right)^{2}\right]\,.$ (29)
We now find the cancellation condition for the soft-scale breaking term, $M^{2}(\sigma_{0})\equiv 0$, to be
$(yv_{\Phi})\sim 262\,{\rm MeV}\,$, (30)
or equivalently,
$\frac{y}{\sqrt{\lambda}}\sim 1.3\times\left(\frac{100\,{\rm MeV}}{\mu_{5}}\right)\,$, (31)
where we have approximated ${\cal S}\sim\sigma_{0}$ in the log factors. As has been discussed in Zhang:2025kbu , more precisely, the size of the soft-scale breaking term is constrained to realize the CW-type supercooling as
$\Bigg|\frac{M^{2}(\sigma_{0})}{\langle{\sigma_{0}}\rangle^{2}}\Bigg|\lesssim{\cal O}(10^{-10})\,$,
${\rm i.e.,}\qquad$ $\Bigg|1-\left(\frac{yv_{\Phi}}{262\,{\rm MeV}}\right)^{2}\Bigg|\lesssim{\cal O}(10^{-11})\,$, (32)
for $\langle\sigma_{0}\rangle=$ 400 MeV - 700 MeV.